OpenStudy (mathlegend):

What two numbers is radical 52 between?

OpenStudy (mathlegend):

\[\sqrt{52}\]

OpenStudy (mathlegend):

Is it between 7 & 8

OpenStudy (goformit100):

YEAH

OpenStudy (anonymous):

7 and 8

OpenStudy (anonymous):

\[\sqrt{52}= \sqrt{2*2*13} =2\sqrt{13}\]